# On the size of universal graphs for spanning trees

### v1

@@ Abstract
For a graph family $\mathcal{F}$, an $\mathcal{F}$-universal graph is a graph which contains all elements of $\mathcal{F}$ as subgraphs. The minimum number of edges in $\mathcal{F}$-universal graphs has been extensively studied for various graph families $\mathcal{F}$. To list a few, the family of all bounded degree graphs [2, 3, 4, 9], the family of all $m$-edge graphs [1, 5, 7], the families of planar graphs with certain restrictions [6, 8, 15], the families of separable bounded degree graphs [10] and the family of trees [11, 13, 12, 14, 16, 18].
Chung and Graham [11, 12] considered the natural problem of determining the minimum number of edges of a $\mathcal{T}_{n}$-universal graph, where $\mathcal{T}_{n}$ is the family of all $n$-vertex trees. They define $s^{*}(n)$ as the minimum number of edges in such a graph. Additionally, they define $s(n)$ as the minimum number of edges in $n$-vertex $\mathcal{T}_{n}$-universal graphs. In other words, $s(n)$ is the minimum number of edges in an $n$-vertex graph that contains all possible spanning trees.
Chung and Graham [11] proved a lower bound on $s^{*}(n)$ and Chung and Graham [12] claimed an upper bound on $s(n)$ yielding the following bound.
@@
We explain this mistake and how to correct it in Section˜3. We note that this mistake was independently discovered by Kaul and Wood in [17], who provided an alternative weaker bound $s(n)=O(n\log n\log\log n)$. In fact, Bhatt, Chung, Leighton, and Rosenberg [6] and Chung [10] both cited the upper bound from [12] as $\frac{7}{2}n\log_{2}n+O(n)$. So we believe that they knew about the mistake and the way to fix it as we present in Section˜3.
Recently, Győri, Li, Salia and Tompkins in [16] improved the lower bound to $n\ln{n}-O(n)\leq s^{*}(n).$ So the state of the art is
@@
Here, the gap between two bounds are a multiplication factor of a number $\frac{7}{2\ln(2)}=5.04\dots$, slightly larger than $5$. On the other hand, the upper bound has remained unchanged since [12] for over four decades. In this paper, we present the first improvement since [12].
@@ 2 Preliminaries
In this section, we collect some useful definitions and lemmas. Many of the definitions in this section are modifications of the ones from [12].
@@ 2.1 Proof overview
As explained in Section˜3, Chung and Graham [12] considered a graph generated by a complete binary tree $B(k)$ and obtained a universal graph with $\frac{7}{2}n\log_{2}n+O(n)$ edges. One natural idea for further improvement is to consider graphs generated by other complete $t$-ary trees for larger values of $t$. It turns out that complete ternary tree yields a slightly better example with $5n\log_{3}n+O(n)\leq 3.156\cdot n\log_{2}n$ edges while all $t>3$ yield worse examples.
To explain the rough idea in [12], suppose that we want to embed a tree $T^{\prime}$ into $G_{T}$, where $G_{T}$ looks like Figure˜1. Then one can use Lemma˜2.1 to find a vertex $w$ and some $w$-components $C_{1},\dots,C_{t}$ of $T^{\prime}$ such that $C_{[t]}$ has size between $\nu(u)$ and $2\nu(u)$. Take a subgraph $G_{2}$ induced by $D_{T}[l(u)]\cup D_{T}[u]\cup\{u^{*}\}$ as in Figure˜1. By showing this is isomorphic to a $T$-admissible graph, we can use the induction hypothesis to find a $T$-admissible embedding $\phi$ of $T^{\prime}_{0}=T^{\prime}[\{w\}\cup C_{[t]}]$ into $G_{2}$ as desired. Furthermore, we ensure that $w$ maps to $u$. Once again, we apply the induction hypothesis to find an embedding $\phi^{\prime}$ of $T^{\prime}_{1}=T^{\prime}\setminus T^{\prime}_{0}$ into $G_{T}-\phi(T^{\prime}_{0})$. As all the edges between $T^{\prime}_{0}$ and $T^{\prime}_{1}$ are incident to $w$ and $u=\phi(w)$ is adjacent to every vertex in $G_{T}$, we can concatenate $\phi$ and $\phi^{\prime}$ to obtain a desired embedding. One crucial condition for this argument to work is that $T^{\prime}_{0}$ has fewer than $|G_{2}|$ vertices. This condition is implied by $\nu(l(u))\geq\nu(u)$. Thus, this argument requires that $l(u)$ has at least as many descendants as $u$ for every vertex $u$.
@@
We point out an error in [12] and provide a fix for it in Section˜3. Our main lemma (Lemma˜4.1) for universality is proved in Section˜4. In Section˜5, we introduce a specific (2,1)(2,1)–tree $T$ that satisfies the conditions in Lemma˜4.1. Utilizing the imbalance of $\nu(l(u))$ and $\nu(u)$ in this tree and further structures of $T$, we prove in Section˜5 that the graph $G^{2}_{T}$ has at most $\frac{14}{3}n\log_{3}n+O(n)$ edges as desired. In Section˜6, we introduce another parameter $s^{int}(n)$ as a variation of $s(n)$ and provide a lower bound for it.
@@
3 A remark on the proof in [12].
In this section, we discuss the mistake in [12]. In order to do this, we will first briefly introduce the definitions from [12]. Let $B(k)$ be a perfect binary tree of level $k$, and $V(k)$ be the vertex set of $B(k)$. Again, $B(k)$-admissible set is defined as before.
@@
Note that (iii) is different from $G_{B(k)}$ as defined in Definition˜2.1. Here, $w$ is the nearest-left sibling of the parent of $u$, instead of the nearest-left cousin. In fact, if the parent of $u$ is the left-most child of its parent, (iii) does not add any more edges in Definition˜3.1 while Definition˜2.1 does. We say that a graph $H$ is $G(k)$-admissible if it is an induced subgraph $G(k)[S]$ of $G(k)$ for some $B(k)$-admissible set $S$. The following is the main lemma from [12].
@@
(Lemma labeled (*) in [12]) Let $S$ be a $B(k)$-admissible set and $G=G(k)[S]$ be the subgraph of $G(k)$ induced by $S$. Let $T$ be any tree with at most $|S|$ vertices and $v\in V(T)$. Then there exists an embedding $\lambda:T\rightarrow G$ such that the vertex set $V(G)-\lambda(V(T))$ is $B(k)$-admissible. Furthermore, the location of $\lambda(v)$ could be determined as follows:
@@
See Figure˜1 (which is same as Figure 7 in [12]). It explains the part II, case (ii), subcase (b) of the proof for Lemma˜3.1 in [12]. It was claimed in [12] that the graph $G_{2}$ in Figure˜1 is isomorphic to an $G(K)$-admissible graph, but this is not true.
@@ 2. $\Phi$2
$\phi(w)=v_{1}\enspace\text{ and }\enspace\phi(x_{1})=\left\{\begin{array}[]{ll}v_{2}&\text{ if }n^{\prime}=n-1.\\ v&\text{ if }n^{\prime}=n.\end{array}\right$.
@@
$x=\nu(v_{t}),\enspace y=\nu(v_{t-1})\enspace\text{and}\enspace z=\nu(v_{t-2}).$
@@
$T^{\prime}_{0}=T^{\prime}[\{w\}\cup C_{[p]}]\enspace\text{ and }\enspace T^{\prime}_{1}=T^{\prime}-T^{\prime}_{0}.$
@@
$T^{\prime\prime}_{0}=T^{\prime}[C^{\prime}\cup C_{0}],\enspace T^{\prime\prime}_{1}=T^{\prime}_{0}[C_{1}]-C^{\prime}\enspace\text{ and }\enspace T^{\prime\prime}_{2}=T^{\prime}[C_{2}].$
@@ 2. $\bullet$
$a_{\ell+1,1},\enspace a_{\ell+1,1}+a_{\ell+1,2}+1,\enspace a_{\ell+1,1},\enspace a_{\ell+1,1}+a_{\ell+1,2}+1,\enspace 2a_{\ell+1,1}+a_{\ell+1,2}+2,\enspace 2a_{\ell+1,1}+2a_{\ell+1,2}+3$

### v2

@@ Abstract
For a graph family $\mathcal{F}$, an $\mathcal{F}$-universal graph is a graph which contains all elements of $\mathcal{F}$ as subgraphs. The minimum number of edges in $\mathcal{F}$-universal graphs has been extensively studied for various graph families $\mathcal{F}$. To list a few, the family of all bounded degree graphs [alonsecond, szemeredi, alonthird, capalbosecond], the family of all $m$-edge graphs [alon, babai, bucic], the families of planar graphs with certain restrictions [bhatt, capalbo, joret], the families of separable bounded degree graphs [chung] and the family of trees [grahamfirst, grahamsecond, grahamthird, pipenger, gyori, montgomery].
Chung and Graham [grahamfirst, grahamthird] considered the natural problem of determining the minimum number of edges of a $\mathcal{T}_{n}$-universal graph, where $\mathcal{T}_{n}$ is the family of all $n$-vertex trees. They define $s^{*}(n)$ as the minimum number of edges in such a graph. Additionally, they define $s(n)$ as the minimum number of edges in $n$-vertex $\mathcal{T}_{n}$-universal graphs. In other words, $s(n)$ is the minimum number of edges in an $n$-vertex graph that contains all possible spanning trees.
Chung and Graham [grahamfirst] proved a lower bound on $s^{*}(n)$ and Chung and Graham [grahamthird] claimed an upper bound on $s(n)$ yielding the following bound.
@@
We explain this mistake and how to correct it in Section˜3. We note that this mistake was independently discovered by Kaul and Wood in [kaul], who provided an alternative weaker bound $s(n)=O(n\log n\log\log n)$. In fact, Bhatt, Chung, Leighton, and Rosenberg [bhatt] and Chung [chung] both cited the upper bound from [grahamthird] as $\frac{7}{2}n\log_{2}n+O(n)$. So we believe that they knew about the mistake and the way to fix it as we present in Section˜3.
Recently, Győri, Li, Salia and Tompkins in [gyori] improved the lower bound to $n\ln{n}-O(n)\leq s^{*}(n).$ So the state of the art is
@@
Here, the gap between two bounds are a multiplication factor of a number $\frac{7}{2\ln(2)}=5.04\dots$, slightly larger than $5$. On the other hand, the upper bound has remained unchanged since [grahamthird] for over four decades. In this paper, we present the first improvement since [grahamthird].
@@ 2 Preliminaries
In this section, we collect some useful definitions and lemmas. Many of the definitions in this section are modifications of the ones from [grahamthird].
@@ 2.1 Proof overview
As explained in Section˜3, Chung and Graham [grahamthird] considered a graph generated by a complete binary tree $B(k)$ and obtained a universal graph with $\frac{7}{2}n\log_{2}n+O(n)$ edges. One natural idea for further improvement is to consider graphs generated by other complete $t$-ary trees for larger values of $t$. It turns out that complete ternary tree yields a slightly better example with $5n\log_{3}n+O(n)\leq 3.156\cdot n\log_{2}n$ edges while all $t>3$ yield worse examples.
To explain the rough idea in [grahamthird], suppose that we want to embed a tree $T^{\prime}$ into $G_{T}$, where $G_{T}$ looks like Figure˜1. Then one can use Lemma˜2.1 to find a vertex $w$ and some $w$-components $C_{1},\dots,C_{t}$ of $T^{\prime}$ such that $C_{[t]}$ has size between $\nu(u)$ and $2\nu(u)$. Take a subgraph $G_{2}$ induced by $D_{T}[l(u)]\cup D_{T}[u]\cup\{u^{*}\}$ as in Figure˜1. By showing this is isomorphic to a $T$-admissible graph, we can use the induction hypothesis to find a $T$-admissible embedding $\phi$ of $T^{\prime}_{0}=T^{\prime}[\{w\}\cup C_{[t]}]$ into $G_{2}$ as desired. Furthermore, we ensure that $w$ maps to $u$. Once again, we apply the induction hypothesis to find an embedding $\phi^{\prime}$ of $T^{\prime}_{1}=T^{\prime}\setminus T^{\prime}_{0}$ into $G_{T}-\phi(T^{\prime}_{0})$. As all the edges between $T^{\prime}_{0}$ and $T^{\prime}_{1}$ are incident to $w$ and $u=\phi(w)$ is adjacent to every vertex in $G_{T}$, we can concatenate $\phi$ and $\phi^{\prime}$ to obtain a desired embedding. One crucial condition for this argument to work is that $T^{\prime}_{0}$ has fewer than $|G_{2}|$ vertices. This condition is implied by $\nu(l(u))\geq\nu(u)$. Thus, this argument requires that $l(u)$ has at least as many descendants as $u$ for every vertex $u$.
@@
We point out an error in [grahamthird] and provide a fix for it in Section˜3. Our main lemma (Lemma˜4.1) for universality is proved in Section˜4. In Section˜5, we introduce a specific (2,1)(2,1)–tree $T$ that satisfies the conditions in Lemma˜4.1. Utilizing the imbalance of $\nu(l(u))$ and $\nu(u)$ in this tree and further structures of $T$, we prove in Section˜5 that the graph $G^{2}_{T}$ has at most $\frac{14}{3}n\log_{3}n+O(n)$ edges as desired. In Section˜6, we introduce another parameter $s^{int}(n)$ as a variation of $s(n)$ and provide a lower bound for it.
@@
3 A remark on the proof in [grahamthird].
In this section, we discuss the mistake in [grahamthird]. In order to do this, we will first briefly introduce the definitions from [grahamthird]. Let $B(k)$ be a perfect binary tree of level $k$, and $V(k)$ be the vertex set of $B(k)$. Again, $B(k)$-admissible set is defined as before.
@@
Note that (iii) is different from $G_{B(k)}$ as defined in Definition˜2.1. Here, $w$ is the nearest-left sibling of the parent of $u$, instead of the nearest-left cousin. In fact, if the parent of $u$ is the left-most child of its parent, (iii) does not add any more edges in Definition˜3.1 while Definition˜2.1 does. We say that a graph $H$ is $G(k)$-admissible if it is an induced subgraph $G(k)[S]$ of $G(k)$ for some $B(k)$-admissible set $S$. The following is the main lemma from [grahamthird].
@@
(Lemma labeled (*) in [grahamthird]) Let $S$ be a $B(k)$-admissible set and $G=G(k)[S]$ be the subgraph of $G(k)$ induced by $S$. Let $T$ be any tree with at most $|S|$ vertices and $v\in V(T)$. Then there exists an embedding $\lambda:T\rightarrow G$ such that the vertex set $V(G)-\lambda(V(T))$ is $B(k)$-admissible. Furthermore, the location of $\lambda(v)$ could be determined as follows:
@@
See Figure˜1 (which is same as Figure 7 in [grahamthird]). It explains the part II, case (ii), subcase (b) of the proof for Lemma˜3.1 in [grahamthird]. It was claimed in [grahamthird] that the graph $G_{2}$ in Figure˜1 is isomorphic to an $G(K)$-admissible graph, but this is not true.
@@ 2. $\Phi$2
$\phi(w)=v_{1}\kern 5.0pt\text{ and }\kern 5.0pt\phi(x_{1})=\left\{\begin{array}[]{ll}v_{2}&\text{ if }n^{\prime}=n-1.\\ v&\text{ if }n^{\prime}=n.\end{array}\right$.
@@
$x=\nu(v_{t}),\kern 5.0pty=\nu(v_{t-1})\kern 5.0pt\text{and}\kern 5.0ptz=\nu(v_{t-2}).$
@@
$T^{\prime}_{0}=T^{\prime}[\{w\}\cup C_{[p]}]\kern 5.0pt\text{ and }\kern 5.0ptT^{\prime}_{1}=T^{\prime}-T^{\prime}_{0}.$
@@
$T^{\prime\prime}_{0}=T^{\prime}[C^{\prime}\cup C_{0}],\kern 5.0ptT^{\prime\prime}_{1}=T^{\prime}_{0}[C_{1}]-C^{\prime}\kern 5.0pt\text{ and }\kern 5.0ptT^{\prime\prime}_{2}=T^{\prime}[C_{2}].$
@@ 2. $\bullet$
$a_{\ell+1,1},\kern 5.0pta_{\ell+1,1}+a_{\ell+1,2}+1,\kern 5.0pta_{\ell+1,1},\kern 5.0pta_{\ell+1,1}+a_{\ell+1,2}+1,\kern 5.0pt2a_{\ell+1,1}+a_{\ell+1,2}+2,\kern 5.0pt2a_{\ell+1,1}+2a_{\ell+1,2}+3$